60 1/4 -58 1/18 need to find the height of the man over his son. I have done this much 1/4=8/8=8/32 then 1/18 =2/2 = 2/32 then I did 60 8/32- 58 2/32 = 2 4/32 for the differance this is where I get stuck how do I break that down to smaller
Found 2 solutions by Fombitz, stanbon:
Answer by Fombitz(32388)
Answer by stanbon(75887)

60 1/4 -58 1/18 need to find the height of the man over his son. I have done this much 1/4*8/8=8/32 then 1/18*2/2 = 2/36 then I did 60 8/32- 58 2/36
= 2 + [1/4 - 1/18] = 2 + [(18-4)/(4*18)] = 2 + [14/
------
60 1/4 -58 1/18
----
= (60-58) + [(1/4)-(1/18)]
-----
= 2 + [(18-4)/(4*18)]
-----
= 2 + (14/(4*18)
-----
= 2 + (7/36)
-----
= 2.1944
================
========
Cheers,
Stan H.